Silicon carbide, SiC, is made by heating solid carbon in the presence of silicon dioxide gas, (SiO2). Carbon monoxide is the side product of the reaction.


Calculate the mass(in gram) of SiC formed when 3.00 g of SiO2 is allowed to react with 4.50 g of carbon.


1
Expert's answer
2021-09-06T02:34:59-0400

2C + SiO2 = SiC + 2CO

n (SiC) = n (SiO2) = 2 x n (C)

n = m / M

M (C) = 12 g/mol

M (SiO2) = 60.08 g/mol

M (SiC) = 40.11 g/mol

n (C) = 4.50 / 12 = 0.375 mol

n (SiO2) = 3.0 / 60.08 = 0.05 mol

SiO2 is the limiting reactant. Therefore:

n (SiC) = 2 x n (SiO2) = 2 x 0.05 = 0.1 mol

m (SiC) = n x M = 0.1 x 40.11 = 4.01 g